# Quantifying the Effect of Test Set Contamination on Generative Evaluations

### v1

@@ Abstract
As frontier AI systems are pretrained on web-scale data, test set contamination has become a critical concern for accurately assessing their capabilities. While research has thoroughly investigated the impact of test set contamination on discriminative evaluations like multiple-choice question-answering, comparatively little research has studied the impact of test set contamination on generative evaluations. In this work, we quantitatively assess the effect of test set contamination on generative evaluations through the language model lifecycle. We pretrain language models on mixtures of web data and the MATH benchmark, sweeping model sizes and number of test set replicas contaminating the pretraining corpus; performance improves with contamination and model size. Using scaling laws, we make a surprising discovery: including even a single test set replica enables models to achieve lower loss than the irreducible error of training on the uncontaminated corpus. We then study further training: overtraining with fresh data reduces the effects of contamination, whereas supervised finetuning on the training set can either increase or decrease performance on test data, depending on the amount of pretraining contamination. Finally, at inference, we identify factors that modulate memorization: high sampling temperatures mitigate contamination effects, and longer solutions are exponentially more difficult to memorize than shorter ones, presenting a contrast with discriminative evaluations, where solutions are only a few tokens in length. By characterizing how generation and memorization interact, we highlight a new layer of complexity for trustworthy evaluation of AI systems.
@@ 1 Introduction
Test set contamination – the inclusion of benchmarks in pretraining data – has emerged as a critical threat to the trustworthy evaluation of frontier AI systems (Sainz et al., 2023; Schaeffer, 2023; Xu et al., 2024a; Deng et al., 2024a; Reuel et al., 2025). Evaluation aims to measure generalization on tasks the model has never seen, yet the sheer scale of modern pretraining makes such contamination increasingly likely (Brown et al., 2020; Du et al., 2022; Wei et al., 2022a; Chowdhery et al., 2022; Touvron et al., 2023).
Prior research has sought to quantify the impact of contamination through two primary lenses. Statistical approaches attempt to detect contamination or estimate its influence by by modifying the test set, for example, by reordering, rephrasing, or replicating benchmark problems, e.g., (Oren et al., 2023; Ni et al., 2025; Shi et al., 2024; Golchin & Surdeanu, 2023, 2024; Roberts et al., 2024; Wang et al., 2025; Zhang et al., 2024a). In comparison, controlled approaches – which offer the most rigorous measurement – intentionally contaminate pretraining corpora to quantify how specific dosages of leakage translate into inflated performance, e.g., (Magar & Schwartz, 2022; Jiang et al., 2024; Oren et al., 2023; Yao et al., 2024; Wang et al., 2025; Kocyigit et al., 2025; Bordt et al., 2025). For a more thorough discussion, please see Appendix A Related Work.
@@
Figure 1: Performance on Generative Benchmarks Increases with Test Set Contamination and Model Size. Schematic: We pretrained compute-optimal language models (3434M–344344M parameters) on corpora containing different replicas of the MATH test set (031623162). Evaluation used greedy decoding (temperature =0=0). Left: Math Verify scores (top) rise and cross entropy on the test set (bottom) falls as contamination increases, consistent with discriminative evaluations, with a sharp improvement around evaluations, with a sharp improvement around $100$ replicas. Right: The ratio in the loss on the MATH test set between $R$ replicas and 0 grows with model size, meaning larger models benefit more from test set contamination for the same number of replicas.
@@
However, with the rapid advancement of model capabilities and the advent of reasoning models (OpenAI et al., 2024a; Google Gemini Team et al., 2025; Xu et al., 2025), the field is shifting towards generative benchmarks, or benchmarks for which the model must generate an answer rather than choosing between provided answers. Whether test set contamination has the same effect on generative evaluation as on discriminative evaluations is unclear. Discriminative evaluations typically require the model to place higher probability mass on the correct choice than on a small number of alternative incorrect choices (Gao et al., 2024; Schaeffer et al., 2025d), and candidate choices are often only a couple of tokens long. In contrast, generative evaluations require the model to produce solutions spanning tens-to-thousands of tokens without decohering, and introduce new considerations such as the sampling temperature (Ackley et al., 1985), the sampling algorithm (e.g., top-k (Fan et al., 2018), top-p (Holtzman et al., 2020)) and the solution length.
In this work, we quantitively study the effects of test set contamination on generative reasoning, focusing on the widely used MATH benchmark (Hendrycks et al., 2021b). We pretrain dozens of language models on corpora contaminated with varying numbers of test set replicas, sweeping across model sizes, sampling temperatures, and token budgets. Our findings reveal that while generative contamination shares some similarities with discriminative settings, it introduces new interactions between memorization and generalization.
@@
In this work, we make the following contributions:
@@
  • Pretraining (Scaling & Irreducible Error): We quantify how contamination impacts the initial pretraining phase. We find that performance scales with contamination frequency and model size, similar to discriminative settings. Using neural scaling laws, we discovered that including even a single replica of the test set enables models to achieve a lower loss than the irreducible error achievable with the uncontaminated corpus.
@@
  • Further Training (Overtraining & Supervised Finetuning): We find that training beyond compute-optimal with fresh data dilutes the effects of contamination, similar to discriminative settings. We then show that SFT on the training data has opposing effects, depending on the amount of contamination in pretraining: performance improves for low contamination, but worsens for high contamination.
  • Inference (Temperature & Solution Length): We identify distinct factors that modulate memorization during generation: increasing sampling temperature and increasing solution sequence length reduces model performance, as models struggle to regurgitate long sequences without decohering.
@@
  • Correction to Evaluation Library: We identify and fix a critical implementation error in the widely used EleutherAI LM Evaluation Harness (Gao et al., 2024) for Math Verify Scores. Our correction ensures that valid reference solutions are accurately scored as correct (raising gold reference solutions’ scores from $\mathord{\sim}$70% to 100%), a necessary step for trustworthy reporting on mathematical benchmarks.
2 Methodology
@@
Pretraining To study the effect of contamination in generative evaluations, we pretrained transformer (Vaswani et al., 2017) causal language models from initialization using the Qwen 3 architecture (Yang et al., 2025a), sweeping model sizes: 3434M, 6262M, 9393M, 153153M, 344344M. Following Chinchilla compute-optimal scaling (Hoffmann et al., 2022), each model was pretrained with 20 tokens-per-parameter. For each model size and token budget, we created multiple pretraining corpora by contaminating a high quality web crawl corpus (Penedo et al., 2024) with a different number of replicas of the benchmark test set: from 0 (uncontaminated) through $1,3,10,32,100,316,1000,3162$ (uniformly spaced on a log scale). Pretraining compute was calculated using the common approximation $C\approx 6\,N\,D$ (Kaplan et al., 2020; Sardana et al., 2024; Porian et al., 2024; Gadre et al., 2024), where $N$ is model parameters and $D$ is pretraining tokens. For implementation details, see Appendix B.
@@
Benchmark We chose the ubiquitous MATH (Hendrycks et al., 2021b) benchmark of competition math problems. The MATH dataset has several properties that made it our benchmark of choice: it is comparatively large, the answers are easy to verify, and the benchmark includes solutions as well as answers. These solutions exhibit high variability in terms of both length and difficulty. The MATH test set contains $\mathord{\sim}1.4$M tokens under the Qwen 3 tokenizer.
@@
Evaluating We evaluated our models using two metrics. The first metric we report is Math Verify, defined as the fraction of problems for which the model generates solutions that are verified to be mathematically equivalent to the benchmark’s boxed answers. We initially evaluated our models using EleutherAI’s Language Model Evaluation Harness (Gao et al., 2024), but discovered an error with how Math Verify scores are computed; for example, the benchmark’s gold reference solutions obtained a Math Verify score around 70%. We worked with its developers to correct the implementation. As an aside, this suggests to us that any research reporting MATH scores from the past 1+ years may have reported incorrect scores. The second metric we report is the Cross Entropy of the gold reference solutions given the problems, which were previously demonstrated to be useful for studying scaling properties of generative evaluations during pretraining (Schaeffer et al., 2025a). We used temperature-only sampling, beginning with temperature 0 (“greedy”), and expanding to more temperatures in Sec. 5.
@@ 3 Pretraining: Scaling & Irreducible Error
Figure 3: Overtraining with Fresh Data Mitigates Contamination. We find an interaction between contamination and overtraining (i.e., training longer than Chinchilla compute-optimal; Eqn. 2) on fresh data. For models with low contamination, cross entropy on the MATH test set decreases with increasing overtraining; however, for models with high contamination, cross entropy increases with overtraining. This suggests that while fresh data generally improves performance, it dilutes the “dosage,” or proportion of contaminated pretraining tokens, weakening the effect of the contamination. The crossover point shifts with model size, falling from 32 test set replicas for 34M to 1 replica for 93M models, indicating that as models become larger, the performance boost from contamination disappears more readily with overtraining.
@@
Consistent with discriminative evaluations, increasing the number of benchmark replicas in the pretraining corpus increases Math Verify scores and decreases cross entropies (Fig 1 Left), as does increasing the model size (Fig 1 Right). We observe a non-linear relationship between the number of test set replicas and model performance: For low levels of contamination ($\leq$ 10 replicas), the impact on performance is minimal, with Math Verify scores and cross entropies remaining close to uncontaminated performance; at around 100 replicas, performance sharply increases (Fig. 5). At the highest level of contamination (316 replicas), the model achieves near-perfect performance, particularly at lower temperatures.
@@
Finding #2: Scaling Laws Suggest Including A Single Test Set Replica Achieves Lower Loss Than the Irreducible Error of the Uncontaminated Corpus
@@
A natural question to ask is: how much does test set contamination “buy” the model in terms of performance. More specifically, how much pretraining compute must be spent on an uncontaminated pretraining corpus to match the performance of a model trained on a corpus containing $R$ replicas of the benchmark test set?
@@
To answer this question, we turned to neural scaling laws. Based on previous work (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI et al., 2024b; Hu et al., 2024; Schaeffer et al., 2025a), for each scaling series pretrained on corpora contaminated with $R\in\{0,1,3,10,32,100,316\}$ replicas of the MATH test set, we fit neural scaling laws for the cross entropy loss $\mathcal{L}$ on the benchmark test set:
$\mathcal{L}(C,R)=E_{0}(R)+\frac{C_{0}(R)}{C^{\alpha(R)}},$ (1)
where $E_{0}(R)>0$ is the irreducible error, $C_{0}(R)>0$ is the compute prefactor and $\alpha(R)>0$ is the compute exponent.
@@
Fig. 2 shows each scaling law, as well as each scaling law’s estimated parameters as a function of the number of test set replicas $R$. Our models’ pretraining compute budgets and losses on the test set were reasonably well fit by Eqn. 1, and the pretraining prefactors and pretraining exponents were roughly constant for the various values of $R$. The biggest effect of increasing contamination is that the irreducible error shrinks from $E(R=0)=0.359$ to $E(R=312)=0.0347$. On the models we trained, we found that including even a single replica enables almost all models to achieve lower cross entropy losses than the irreducible error achievable using unlimited pretraining compute on an uncontaminated pretraining corpus. Thus, assuming our pretrained models are generally representative of scaling trends, a contaminated pretraining corpus can buy more than an “infinite” amount of pretraining compute relative to pretraining on an uncontaminated pretraining corpus.
Figure 4: Supervised Finetuning on the Train Set Has Opposing Effects, Depending on Pretraining Contamination. For models pretrained with little-to-no contamination (<10<10 test set replicas), supervised finetuning (SFT) on the MATH train set decreases loss on the test set. For models pretrained with more contamination (>10>10 test set replicas), SFTing on the train set increases loss on the test set. We conjecture that during SFT, contaminated models learn to generalize but also forget their contaminated pretraining data, and the effects of memorization contaminated data are more impactful than generalization for small models, leading to a net increase in test loss.
@@
4 Further Training: Overtraining & Supervised Finetuning (SFT)
Finding #3: Overtraining with Fresh Data Mitigates Contamination
Bordt et al. (2025) recently reported that for discriminative evaluations, the effect of contamination diminishes when models are trained beyond the “compute optimal” ratio of 20 tokens-per-parameter (Hoffmann et al., 2022) on fresh data. We tested whether this so-called overtraining (Touvron et al., 2023; Sardana et al., 2024; Gadre et al., 2024; Schaeffer et al., 2025a) has similar effects for generative benchmarks. We extended our Sec. 2 pretraining sweep into the overtrained regime, pretraining on

### v2

@@ Abstract
Test set contamination – the inclusion of benchmarks in pretraining data – is a critical threat to the trustworthy evaluation of frontier AI systems. While its impact on discriminative evaluations is well-studied, contamination on generative evaluations remains underexplored. We quantitatively assess these effects across the language model lifecycle by pretraining models (up to 344M parameters) on web data contaminated with varying numbers MATH test set replicas. While performance expectedly improves with contamination and model size, our scaling law analysis reveals a fundamental breach: including even a single test set replica enables models to achieve lower loss than the irreducible error of training on the uncontaminated corpus. We then study additional training: overtraining with fresh data dilutes contamination effects, whereas supervised finetuning on the training set improves performance for low contamination but degrades it for high contamination. At inference, we identify three distinct regimes of memorization—ranging from exponential decoherence to deterministic lock-in—governed by solution length and sampling temperature. Finally, we identify and fix a critical implementation error in a widely used evaluation library that previously underreported mathematical reasoning performance. By characterizing how generation and memorization interact, we highlight new considerations for trustworthy AI evaluation.
@@ 1 Introduction
As frontier AI systems are pretrained on web-scale data, test set contamination has become a critical concern for accurately assessing their capabilities (Sainz et al., 2023; Schaeffer, 2023; Xu et al., 2024a; Deng et al., 2024a; Reuel et al., 2025). Evaluation aims to measure generalization on tasks the model has never seen, yet the sheer scale of modern pretraining makes such contamination likely (Brown et al., 2020; Du et al., 2022; Wei et al., 2022a; Chowdhery et al., 2022; Touvron et al., 2023; Penedo et al., 2024).
Prior research has sought to quantify the impact of test set contamination, also known as leakage, through two primary lenses. Statistical approaches attempt to detect contamination or estimate its influence by modifying the test set, for example, by reordering, rephrasing, or replicating benchmark problems, e.g., (Oren et al., 2023; Ni et al., 2025; Shi et al., 2024; Golchin & Surdeanu, 2023, 2024; Roberts et al., 2024; Wang et al., 2025; Zhang et al., 2024a). In comparison, controlled approaches – which offer the most rigorous measurement – intentionally contaminate pretraining corpora to quantify how specific doses of leakage inflate performance, e.g., (Magar & Schwartz, 2022; Jiang et al., 2024; Oren et al., 2023; Yao et al., 2024; Duan et al., 2024; Wang et al., 2025; Kocyigit et al., 2025; Bordt et al., 2025; Wei et al., 2025). For a more in-depth discussion, please see Appendix A Related Work.
@@
Figure 1: Performance on Generative Benchmarks Increases with Test Set Contamination and Model Size. Schematic: We pretrained compute-optimal language models (3434M–344344M parameters) on corpora containing different replicas of the MATH test set (031623162). Evaluation used greedy decoding (temperature =0=0). Left: As contamination (quantified by the number of test set replicas in the pretraining corpus) increases, Math Verify scores (top) rise and cross entropies on the test set (bottom) fall, consistent with discriminative evaluations, with a sharp improvement around $100$ test set replicas. Right: The ratio in the loss on the MATH test set between $R$ replicas and 0 replicas grows with model size, meaning larger models benefit more from contamination for the same number of replicas.
@@
However, with the rapid advancement of model capabilities and the advent of reasoning models (OpenAI et al., 2024a; Google Gemini Team et al., 2025; Xu et al., 2025), the field is shifting towards benchmarks for which the model must generate an answer rather than choose between provided answers. Whether test set contamination has the same effect on generative evaluation as on discriminative evaluations is unclear. Discriminative evaluations typically require the model to place higher probability mass on the correct choice than on a small number of alternative incorrect choices (Gao et al., 2024; Schaeffer et al., 2025e), and candidate choices are often only a couple of tokens long. In contrast, generative evaluations require the model to produce solutions spanning tens-to-thousands of tokens without straying from the memorized path, and introduce new considerations such as the sampling temperature (Ackley et al., 1985), the sampling algorithm (e.g., top-k (Fan et al., 2018), top-p (Holtzman et al., 2020)) and the solution length.
@@
In this work, we quantitatively study the effects of test set contamination on generative evaluations, focusing on the widely used MATH benchmark (Hendrycks et al., 2021b). We pretrain dozens of language models on corpora contaminated with varying numbers of test set replicas, sweeping across model sizes, sampling temperatures, and token budgets. We make the following contributions:
@@
  • Pretraining (Scaling & Irreducible Error): We quantify how contamination impacts pretrained models. We find that performance increases with contamination and model size, similar to discriminative settings. Under standard scaling law assumptions, we discover a fundamental breach: including even a single replica of the test set enables achieving a lower loss than the estimated irreducible error of the uncontaminated corpus.
@@
  • More Training (Overtraining & Supervised Finetuning): We find that training beyond compute-optimal with fresh data dilutes the performance gains from contamination. We further reveal that Supervised Finetuning (SFT) has opposing effects depending on the pretraining contamination level: it improves performance for low contamination but actively worsens performance for high contamination models.
  • Inference (Temperature & Solution Length): We identify that sampling temperature and solution length modulate memorization. We mathematically describe three distinct regimes: Exponentially Fast Decoherence, Brittle Memorization, and Deterministic Lock-In. We show that high-temperature sampling acts as a ”truth serum,” decoupling contamination gains from robust generalization.
@@
  • Correction to Evaluation Library: We identify and fix a critical implementation error in the widely used EleutherAI LM Evaluation Harness (Gao et al., 2024) for Math Verify Scores (Appendix C). Our correction ensures that valid reference solutions are accurately scored as correct (raising gold reference solutions’ scores from $\mathord{\sim}$70% to 100%), a necessary step for trustworthy reporting on mathematical benchmarks.
Our investigation follows the standard language model lifecycle: we first examine contamination during pretraining (Sec. 3), then study how subsequent training—both overtraining and supervised finetuning—interacts with contamination (Sec. 4), and finally analyze how inference-time parameters like sampling temperature and output length modulate performance during deployment (Sec. 5).
2 Methods
@@
Pretraining To study the effects of contamination in generative evaluations, we pretrained transformer-based causal language models (Vaswani et al., 2017; Brown et al., 2020) using the Qwen 3 architecture (Yang et al., 2025a), sweeping model sizes: 3434M, 6262M, 9393M, 153153M, 344344M. Each model was pretrained with 20 tokens-per-parameter (Hoffmann et al., 2022). For each model size, we created multiple pretraining corpora by contaminating a high quality web crawl corpus (Penedo et al., 2024) with a different number of replicas of the benchmark test set: from 0 (uncontaminated) through $1,3,10,32,100,316,1000,3162$ (logarithmically uniform). Pretraining compute (floating point operations; FLOP), was approximated as $C\approx 6\,N\,D$ (Kaplan et al., 2020; Sardana et al., 2024; Porian et al., 2024; Gadre et al., 2024), where $N$ is model parameters and $D$ is pretraining tokens. We trained one model for each pair of model size and number of test set replicas. For more details, see App. B.
@@
Benchmark We chose the ubiquitous MATH (Hendrycks et al., 2021b) benchmark of competition math problems due to several properties: it is comparatively large (5,000 test problems), answers exist, and the benchmark includes solutions as well as answers. The MATH test set contains $\mathord{\sim}1.4$M tokens under the Qwen 3 tokenizer.
@@
Evaluation We evaluated our models using two metrics. The first metric we report is Math Verify (Kydlicek et al., 2025), defined as the fraction of problems for which the model generates solutions that are verified to be mathematically equivalent to the benchmark’s boxed answers. We initially evaluated our models using EleutherAI’s Language Model Evaluation Harness (Gao et al., 2024), but discovered a bug in how Math Verify scores are computed by the Harness; for example, the benchmark’s gold reference solutions obtained a Math Verify score of only $\mathord{\sim}70\%$. We worked with the developers to correct the implementation (see Appendix C for details). This suggests that research reporting math_minerva scores with the Eval Harness prior to task version 3.0 may have reported incorrect scores. The second metric we report is the Cross Entropy of the gold reference solutions given the problems, which were previously demonstrated to be useful for studying scaling properties of generative evaluations during pretraining (Schaeffer et al., 2025b). We used temperature-only sampling (Schaeffer et al., 2025a), beginning with temperature 0 (“greedy”), and expanding to more temperatures in Sec. 5.
@@ 3 Pretraining: Scaling & Irreducible Error
We begin at the first stage of the model lifecycle: pretraining. We establish foundational results on how test set contamination interacts with model scale and compute, and use scaling laws to quantify what contamination buys a model.
@@
Consistent with discriminative evaluations, increasing the number of benchmark replicas in the pretraining corpus increases Math Verify scores and decreases cross entropies (Fig. 1 Left), as does increasing model size (Fig. 1 Right). We observe a non-linear relationship between the number of test set replicas and model performance: For low levels of contamination ($\leq$ 10 replicas), the impact on performance is minimal, with Math Verify scores and cross entropies remaining close to uncontaminated performance; at around 100 replicas, performance sharply increases (Fig. 5). At the highest levels of contamination, the model achieves ceiling performance.
Finding #2: Contamination-Driven Performance Is Not Generalization
To determine if the performance gains stemmed from robust generalization or superficial memorization, we evaluated our contaminated models on two modified versions of the MATH test set. First, we rephrased the problems, retaining the original numerical values and logic but altering the linguistic surface form. Second, we perturbed the problems, modifying the numerical values and correct answers while maintaining the problem structure.
In both conditions, performance regresses to match the uncontaminated model across all model sizes and contamination levels in both conditions (Tab. 1). This strongly suggests that the capabilities gained from test set contamination rely almost exclusively on verbatim memorization of the specific test sequences, rather than the acquisition of underlying mathematical reasoning.
@@
Finding #3: Scaling Laws Suggest Including A Single Test Set Replica Achieves Lower Loss Than the Irreducible Error of the Uncontaminated Corpus
Model Replicas Rephrased Perturbed
344M 0 0.04% 0.00%
1 0.00% 0.00%
3 0.00% 0.00%
10 0.00% 0.00%
32 0.00% 0.00%
100 0.02% 0.00%
316 0.00% 0.00%
1000 0.00% 0.00%
3162 0.04% 0.00%
Table 1: Contamination-Driven Performance Is Not Generalization. When MATH test problems are rephrased (same numbers, different wording) or perturbed (same wording, different numbers), model performance collapses to baseline regardless of contamination level or model size, confirming that performance gains from contamination are not generalized mathematical reasoning. Results were consistent across all model sizes.
@@
How much does test set contamination “buy” the model in terms of performance? More specifically, how much pretraining compute must be spent on an uncontaminated pretraining corpus to match the performance of a model trained on a corpus containing $R$ replicas of the benchmark test set?
@@
To answer this question, we turned to neural scaling laws. Based on previous work (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI et al., 2024b; Hu et al., 2024; Schaeffer et al., 2025b), for each scaling series pretrained on corpora contaminated with $R\in\{0,1,3,10,32,100,316\}$ replicas of the MATH test set, we fit neural scaling laws for the cross entropy loss $\mathcal{L}$ on the benchmark test set as a function of pretraining compute $C$, measured in 6ND6\,N\,D FLOP:
$\mathcal{L}(C,R)=E_{0}(R)+C_{0}(R)\cdot C^{-\alpha(R)}.$ (1)
@@
Fig. 2 shows each scaling law, as well as each scaling law’s estimated parameters as a function of the number of test set replicas $R$. Our models’ pretraining compute budgets and losses on the test set are reasonably well fit by Eqn. 1, and the pretraining prefactors and pretraining exponents are roughly constant for the various values of $R$. The biggest effect of increasing contamination is that the irreducible error shrinks from $E=3.594\rightarrow 0.0347$ as $R=0\rightarrow 316$. On the models we trained, including even a single replica enables almost all models to achieve lower cross entropy losses than the estimated irreducible error of the uncontaminated pretraining corpus. Under standard scaling law assumptions – specifically, that the scaling law extrapolates infinitely – a contaminated pretraining corpus can buy more than an “infinite” amount of pretraining compute relative to pretraining on an uncontaminated pretraining corpus.
@@
Figure 3: Overtraining with Fresh Data Mitigates Contamination. Consistent with discriminative evaluations (Bordt et al., 2025), we find an interaction between contamination and overtraining (i.e., training longer than Chinchilla compute-optimal; Eqn. 2) on new fresh data. For models with low contamination, cross entropy on the MATH test set decreases with increasing overtraining; however, for models with high contamination, cross entropy increases with overtraining. This suggests that while fresh data generally improves performance, it dilutes the “dose,” or proportion of contaminated pretraining tokens, thereby lessening the effect of the contamination. The crossover point shifts with model size, falling from 32 test set replicas for 34M to 1 replica for 93M models, indicating larger models lose their contamination advantage more readily when overtrained.
4 Further Training: Overtraining & Supervised Finetuning
After pretraining, models typically undergo additional training. We now examine how two common additional training interventions—overtraining on fresh data and supervised finetuning—interact with pretraining contamination.
Finding #4: Overtraining with Fresh Data Mitigates Contamination
Bordt et al. (2025) recently reported that for discriminative evaluations, the effect of contamination diminishes when models are trained beyond “compute optimal” on fresh data. We tested whether this so-called overtraining (Touvron et al., 2023; Sardana et al., 2024; Gadre et al., 2024; Schaeffer et al., 2025b) has similar effects for generative benchmarks. We extended our Sec. 2 pretraining sweep into the overtrained regime, pretraining on